Find the nth term of the arithmetic sequence,
a(1 symbol)=8, d=3,n=16

Answer by venugopalramana(3286)

Find the nth term of the arithmetic sequence,
a(1 symbol)=8, d=3,n=16
formula for n th.term of A.P.=Tn=a+(n-1)*d...substituting
Tn=T16=8+(16-1)*3=8+15*3=8+45=53
hence T16=53